1- Find the velocity of VB, Vc, VD, Assume w2=8 rad/s wz @z = 150°C
Answer #1

Solution:

Considering the unit of length of link in centimetre.

The velocity of Point B

Vb=rb\omega

=2*10-2*8

=0.16 m/s

By the rigid body consideration,

Vb=rcb \omegacd

\omegacd= Vb / rcb

=0.16/4*10-2

= 4 rad/s

thus,

Vd= rcd*\omegacd

=5.5*10-2*4

=0.22 m/s

Vc=rcc*\omegacd

=0*4

=0 m/s

Answer: Vb=0.16 m/s , Vc= 0 m/s and Vd= 0.22 m/s